OpenStudy (anonymous):

How to integrate sin2xcos2x/sqrt(1+cos4x)? I know that if I do anti-differentiation, I would get 1+cos4x to have the differential sin2xcos2x, but I can't seem to figure out how 1+cos4x differentiate to get sin2xcos2x

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \sin(2x)\cos(2x) }{ \sqrt{(1+\cos(4x)} } dx\]

OpenStudy (anonymous):

is this ur question ?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\frac{\sin(2x)\cos(2x)}{\sqrt{1+\cos(4x)}}dx }\) You know that, \(\color{#000000 }{ \displaystyle\sin 2z = 2\sin z \cos z }\) (In your case, \(\color{#000000 }{ \displaystyle z:=2x }\)) \(\color{#000000 }{ \displaystyle \int\frac{2\sin(4x)}{\sqrt{1+\cos(4x)}}dx }\)

OpenStudy (solomonzelman):

Then, substitute; \(\color{#000000 }{ \displaystyle p=1+\cos(4x) }\) (Or, you can use any other latter/symbol that you want)

OpenStudy (anonymous):

But sin 2x=2sin 2xcos2x \[2\sin2xcos2x \times \cos2x=(2\sin2x)(2\cos2x)?\] Sorry, but I'm still a bit confused how it would become 2 sin(4x) @SolomonZelman

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sin(2z)=2\sin z \cos z }\) That is the rule, right?

OpenStudy (solomonzelman):

Yes, I made a mistake there

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\frac{\sin(2x)\cos(2x)}{\sqrt{1+\cos(4x)}}dx }\) You know that, \(\color{#000000 }{ \displaystyle\sin 2z = 2\sin z \cos z }\) (In your case, \(\color{#000000 }{ \displaystyle z:=2x }\)) \(\color{#000000 }{ \displaystyle \int\frac{1}{2}\frac{\sin(4x)}{\sqrt{1+\cos(4x)}}dx }\)

OpenStudy (solomonzelman):

The rule is: \(\color{#000000 }{ \displaystyle \sin(2z)=2\sin z \cos z }\) When you substitute \(\color{#000000 }{ \displaystyle 2x }\) instead of \(\color{#000000 }{ \displaystyle z }\) into the equation, you are going to get the following: \(\color{#000000 }{ \displaystyle \sin(2\color{#ff0000}{\cdot 2x})=2\sin \color{#ff0000}{ (2x)} \cos \color{#ff0000}{ (2x)} }\) \(\color{#000000 }{ \displaystyle \sin(4x)=2\sin (2x) \cos (2x) }\) And therefore, \(\color{#000000 }{ \displaystyle (1/2)\cdot \sin(4x)=\sin (2x) \cos (2x) }\)

OpenStudy (solomonzelman):

So, your integral is; \(\color{#000000 }{ \displaystyle \frac{1}{2}\int\frac{\sin(4x)}{\sqrt{1+\cos(4x)}}dx }\) And that can be solved by setting \(\color{#000000 }{ \displaystyle p=1+\cos(4x) }\)

OpenStudy (solomonzelman):

Do you have questions about this?

OpenStudy (anonymous):

@SolomonZelman u are very good with integration and even with explanation :)

OpenStudy (solomonzelman):

thanks ]]

OpenStudy (anonymous):

Thank you!! @SolomonZelman. I get it now. :)

OpenStudy (solomonzelman):

What did you get for the integral?

OpenStudy (solomonzelman):

(what is your answer? if you have an answer now...)

OpenStudy (anonymous):

Although I was just about to ask you. The answer is \[-\frac{ 1 }{ 4 }\sqrt{1+\cos (4x)}+c\]

OpenStudy (anonymous):

And carrying on from where you left off, I arrived at the integral of 1/8 sqrt (1+cos (4x)), because if we consider the 1/2 left off from the previous steps...

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\frac{\sin(2x)\cos(2x)}{\sqrt{1+\cos(4x)}}dx }\) You know that, \(\color{#000000 }{ \displaystyle\sin 2z = 2\sin z \cos z }\) (in your case, \(\color{#000000 }{ \displaystyle z:=2x }\)) \(\color{#000000 }{ \displaystyle \frac{1}{2}\int\frac{\sin(4x)}{\sqrt{1+\cos(4x)}}dx }\) \(\color{#000000 }{ \displaystyle p=1+\cos(4x) }\) \(\color{#000000 }{ \displaystyle dp=-4\sin(4x)dx~~\Longrightarrow~~(-1/4)dp=\sin(4x) dx }\) \(\color{#000000 }{ \displaystyle -\frac{1}{8}\int\frac{1}{\sqrt{~p~}}dp }\) \(\color{#000000 }{ \displaystyle -\frac{1}{8}\int p^{-1/2}dp }\) \(\color{#000000 }{ \displaystyle -\frac{1}{4}p^{1/2} + {\rm C} }\) \(\color{#000000 }{ \displaystyle -\frac{1}{4}\sqrt{1+\cos(4x)}+ {\rm C} }\)

OpenStudy (solomonzelman):

yes, I got the same answer

OpenStudy (anonymous):

Oh, yes. sorry, my careless mistake. I didn't divide by 1/2 (ie multiply by 2) when I was doing the integral of 1+cos(4x). Thank you very much.

OpenStudy (solomonzelman):

I don't really see any mistakes that you made ... all I saw is that you arrived at the same answer as I did....